If 27 is added to a two digit number, the result is a number with the same digits, but in reverse order. The sum of the digits is 11. What is the original number?
Can you please solve this for me step by step?

Answer by stanbon(75887)

If 27 is added to a two digit number, the result is a number with the same digits, but in reverse order. The sum of the digits is 11. What is the original number?
----
Let the 2-digit number be 10t+u.
-----
Equation:
10t+u + 27 = 10u+t
t + u = 11
---------------------
9t-9u = -27
t + u = 11
------------
t - u = -3
t + u = 11
--------------
Add and solve for "t":
2t = 8
t = 4
-----
Solve for "u":
t + u = 4
4 + u = 4
u = 0
=======
The 2-digit number:
10t + u = 40
==================
Cheers,
Stan H.
==================